Give what are asked for in the following problems:


1. The data given below are the weights (lbs) of 15 students.


35 38 39 38 39

45 40 47 38 43

40 42 45 50 48


a) Calculate the mean, median and mode.


b) Interpret the result.


1
Expert's answer
2021-12-14T03:09:15-0500

a)


"35, 38, 38, 38, 39, 39, 40, 40,"

"42, 43, 45, 45, 47, 48, 50"



"n=15"


"mean=\dfrac{1}{n}\displaystyle\sum_{i=1}^{n}x_i=\dfrac{1}{15}(35+38+38+38"

"+39+39+40+40+42+43+45+45"

"+47+48+50)=\dfrac{627}{15}=41.8"

"median=40"

"mode=38, appeared\ 3\ times"

b)


"mode<median<mean"

Right-skewed (Positive Skewness) distribution.